# Reframing the Relationship in Out-of-Distribution Detection

### v1

@@ Abstract
The remarkable achievements of Large Language Models (LLMs) have captivated the attention of both academia and industry, transcending their initial role in dialogue generation. The utilization of LLMs as intermediary agents in various tasks has yielded promising results, sparking a wave of innovation in artificial intelligence. Building on these breakthroughs, we introduce a novel approach that integrates the agent paradigm into the Out-of-distribution (OOD) detection task, aiming to enhance its robustness and adaptability. Our proposed method, Concept Matching with Agent (CMA), employs neutral prompts as agents to augment the CLIP-based OOD detection process. These agents function as dynamic observers and communication hubs, interacting with both In-distribution (ID) labels and data inputs to form vector triangle relationships. This triangular framework offers a more nuanced approach than the traditional binary relationship, allowing for better separation and identification of ID and OOD inputs. Our extensive experimental results showcase the superior performance of CMA over both zero-shot and training-required methods in a diverse array of real-world scenarios.
@@ 1 Introduction
The emergence and development of Large Language Models (LLMs) [8, 84, 7, 2] have significantly reshaped the landscape of Artificial Intelligence (AI), marking a pivotal breakthrough in both academic research and practical applications. These models have not only revolutionized the way we generate conversations but also demonstrated their capacity as intermediary agents with more nuanced roles, facilitating the accomplishment of myriad tasks with unprecedented efficiency and adaptability [72, 75, 83]. The Agent paradigm has been extensively applied across multiple domains and tasks, playing a profound role [57, 58, 25].
An important and challenging task within the field of machine learning is to enhance the robustness of models across diverse scenarios. When an artificial intelligence system encounters data that significantly deviates from its training data distribution, OOD detection becomes crucial for ensuring its reliability and robustness. In the past, most OOD detection methods employed single-modal learning [79, 44, 24]. As CLIP [59] has demonstrated astonishing performance across various downstream tasks, an increasing number of CLIP-based methods for out-of-distribution (OOD) detection have emerged [61, 70, 11, 73, 22].
@@
Figure 2: Comparison between Binary Relationship and Vector Triangle Relationship in OOD detection.
@@
In the realm of OOD detection, the predominant methodology consists of binary segmentation between ID and OOD data, typically facilitated by scoring functions. The effectiveness of this prototype is closely tied to the sophisticated design of these scoring functions. Drawing inspiration from the notable success of the agent-based paradigms, the comparative analysis of ID and OOD data via an agent may present a more refined and comprehensive understanding. Structurally, this paradigm shift from binary segmentation to a vector triangle relationship relational framework holds substantial potential for uncovering deeper insights, potentially transforming the interplay between ID labels and data inputs.
@@
In this paper, we propose the Concept Matching with Agent (CMA) methodology, which integrates neutral textual concept prompts in natural language as Agents within the CLIP-based OOD detection framework. Our framework is illustrated in Figure 1. These agents serve dual roles: as observers and as intermediate hubs that facilitate the interaction between ID Labels and Data inputs. By doing so, we aim to establish a vector triangle relationship among the ID labels, data inputs, and the agents themselves. In this triangular vector relationship (Figure  2), the score of OOD data is diminished due to the collision effect of the Agents, thereby widening the gap between the scores of ID data and OOD data. In other words, images closer to the ID class are less likely to be influenced by neutral text concepts, whereas images from the OOD class are more susceptible to these concepts, resulting in lower scores. The method to achieve our idea is derived from our profound insights into the Language-Vision representation (See Section 3). Building on these insights, we formulate the entire Concept Matching with Agent (CMA) framework.
In summary, our proposed method, CMA, possesses three distinct advantages. (1) Our approach obviates the need for training data, enabling zero-shot OOD detection while leveraging the collision between Agent and both ID and OOD data to effectively widen the gap between the two, ensuring robust performance. This stands in stark contrast to traditional OOD detection methods, which rely on extensive external data for intricate training [79, 44]. (2) Our method exhibits remarkable scalability, allowing for the tailoring of specialized Agents to suit various scenarios, thereby further enhancing performance (See Section 5). This is facilitated by the flexibility of our triangular vector relationship, which can enhance the impact of certain OOD images through specific Agents, thereby reducing their scores. (3) It is noteworthy that the CMA maintains robustness against both hard OOD inputs, encompassing both semantic hard OODs [74] and spurious OODs [50]. This makes our approach a truly practical and viable option. The contributions of our study are summarized as follows:
@@ 2 Preliminaries
Contrastive Vision-Language Models
In comparison to traditional CNN architectures, the ViT [16] leverages the Transformer Encoder framework [68] to accomplish the task of image classification, realizing the possibility of utilizing language model architectures for visual tasks. This provides insights into the field of vision-language representation learning, with CLIP [59] being a notable representative. CLIP employs self-supervised contrastive objectives to embed images and their corresponding textual descriptions into a shared feature space, achieving alignment between the two. Structurally, CLIP, which utilizes a dual-stream architecture, comprises an image encoder $\mathcal{I}:x\to\mathbb{R}^{d}$ and a text encoder $\mathcal{T}:t\to\mathbb{R}^{d}$. After pretraining on a dataset of 400 million text-image pairs, the joint visual-language embedding of CLIP associates objects in various patterns. Due to the robust performance of CLIP, several OOD detection methods based on it have emerged. Nevertheless, the challenge of how to better utilize Language-Vision representation for OOD detection remains a difficult yet significant issue.
Zero-shot OOD Detection
For traditional OOD detection frameworks [24], a common assumption is made of a typical real-world scenario wherein classifier $f$ are trained on ID data categorized as $\mathcal{Y}_{in}=\left\{1,...,C\right\}$ and subsequently deployed in an environment containing samples from unknown classes $y\notin\mathcal{Y}_{in}$, outside the distribution of the ID. The classifier $f$ is then tasked with determining membership to the ID. In contrast, zero-shot OOD detection [48, 51, 21], in line with the current trend of deep learning, leverages pre-trained models on open datasets, eliminating the need for additional training of the model. The approach determines membership to the ID by calculating results through mapping the data into a common space $\mathbb{R}^{d}$. Compared to traditional OOD detection frameworks that necessitate training, the zero-shot OOD detection framework is notably more versatile and practical.
@@ 3 Method
3.1 CLIP-based OOD Detection
For current CLIP-based OOD detection, MCM [48] has become a basic paradigm. Its core idea is to treat text embeddings as "concept prototypes" and evaluate their in-distribution or out-of-distribution properties by measuring the similarity between the input image features and these concept prototypes. Specifically, given a set of ID categories $\mathcal{Y}_{in}$ with a corresponding text description for each category, we first use a pre-trained text encoder $\mathcal{T}$ to convert these text descriptions into $d$ dimensional vectors $\mathbf{c}_{i}=\mathcal{T}(t_{i})\in\mathbb{R}^{d}$, where $i\in\{1,2,...,N\}$ and $N$ is the number of ID categories. For any input image $\mathbf{x}^{\prime}$ whose visual features are extracted by an image encoder $\mathcal{I}$ as $\mathbf{v}^{\prime}=\mathcal{I}(\mathbf{x}^{\prime})\in\mathbb{R}^{d}$, the MCM score is defined as the cosine similarity between the visual features and the closest concept prototype, which is scaled by an appropriate softmax to enhance the separability of ID and OOD samples:
where $\text{sim}(\cdot,\cdot)$ represents cosine similarity, $\hat{y}=\arg\max_{i}\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{i})$ represents the most matching concept index, and $\tau$ is a temperature parameter used to adjust the distribution of similarity.
@@ 3.2 How to construct vector triangle relationships using Agents
Figure 3: Heatmaps depicting the cosine similarity between image inputs and ID concept vectors. In Figure (A), the ID concept vectors consist of sentences containing the word "cat" of varying lengths. It is observed that images tend to align with longer sentences regardless of whether there is a matching ID concept. Concurrently, keywords such as "white" significantly influence image matching. In Figure (B), aside from "cat", no ID concept can be precisely matched with the given images. However, other than cat images, all images exhibit a preference for aligning with a long sentence devoid of tangible objects. Notably, cat image remains unaffected, aligning solely with the ID concept "cat".
@@
  • Different words have different weights in the text description. The more important the words are to the overall description of the picture, the higher the weight.
@@
The causes of these phenomena can be contemplated through the training methods and mechanisms of the CLIP-like models.Given that the common text input for pre-trained models during training is long sentences, they possess higher confidence in inferring long sentences. Additionally, CLIP-liked models undergo contrastive learning training, which enables the model to autonomously focus on the differences between various text descriptions, thereby giving higher weights to key words and skewing the predicted outcomes towards the positive class without being influenced by other words.
Drawing on these significant observations, we propose the inclusion of neutral prompts unrelated to the ID category as Agents. By engaging Agents in a ’collision’ with OOD and ID data, we aim to distance the OOD data from the ID domain.
@@ 3.3 Proposed approach
Based on the aforementioned analysis, Concept Matching with Agent (CMA) employs neutral prompts as Agents to widen the gap between ID and OOD by constructing a vector triangle relationship. Essentially, CMA employs external agents to minimize the maximum score, which is fundamentally distinct from traditional learning methods. Therefore, it does not require additional data or training. Specifically, for a set of textual descriptions $t_{i}$, where $t\in\{1,2,...,N\}$ corresponding to ID categories $\mathcal{Y}_{in}$, we select a certain number of neutral prompts $\mathcal{Y}_{ntc}$ to concatenate with them. The number is generally the same as the number of ID categories $N$, and the concatenated textual concepts $\mathcal{Y}_{con}$ are used as the input of the text encoder $\mathcal{T}$ to output text features $\mathbf{c}_{i}=\mathcal{T}(t_{i})$, where $i\in\{1,2,...,2N\}$, $t_{i}\in\mathcal{Y}_{con}=\{\mathcal{Y}_{in},\mathcal{Y}_{ntc}\}$. Then, it is calculated with the image features output $\mathbf{v}^{\prime}=\mathcal{I}(\mathbf{x}^{\prime})$ by the image encoder $\mathcal{I}$.
@@
where $\text{sim}(\cdot,\cdot)$ represents cosine similarity, $\hat{y}=\mathop{\arg\max}\limits_{1\leq i\leq N}\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{i})$ represents the most matching concept index in ID categories, and $\tau$ is a temperature parameter used to adjust the distribution of similarity.
@@
$g\left(\mathbf{x}^{\prime};\mathcal{Y}_{\mathrm{in}},\mathcal{Y}_{ntc},\mathcal{T},\mathcal{I}\right)=\left\{\begin{array}[]{ll}1,&S_{\mathrm{CMA}}\left(\mathbf{x}^{\prime};\mathcal{Y}_{\text{in }},\mathcal{Y}_{ntc},\mathcal{T},\mathcal{I}\right)\geq\lambda\\ 0,&otherwise\end{array},\right.$

### v2

# Concept Matching with Agent for Out-of-Distribution Detection

@@ Abstract
The remarkable achievements of Large Language Models (LLMs) have captivated the attention of both academia and industry, transcending their initial role in dialogue generation. To expand the usage scenarios of LLM, some works enhance the effectiveness and capabilities of the model by introducing more external information, which is called the agent paradigm. Based on this idea, we propose a new method that integrates the agent paradigm into out-of-distribution (OOD) detection task, aiming to improve its robustness and adaptability. Our proposed method, Concept Matching with Agent (CMA), employs neutral prompts as agents to augment the CLIP-based OOD detection process. These agents function as dynamic observers and communication hubs, interacting with both In-distribution (ID) labels and data inputs to form vector triangle relationships. This triangular framework offers a more nuanced approach than the traditional binary relationship, allowing for better separation and identification of ID and OOD inputs. Our extensive experimental results showcase the superior performance of CMA over both zero-shot and training-required methods in a diverse array of real-world scenarios.
Codehttps://github.com/yuxiaoLeeMarks/CMA
@@ 1 Introduction
The emergence and development of Large Language Models (LLMs) (Chang et al. 2024; Zhao et al. 2023; Brown et al. 2020; Achiam et al. 2023) have significantly reshaped the landscape of Artificial Intelligence (AI), marking a pivotal breakthrough in both academic research and practical applications. These models have not only revolutionized the way we generate conversations but also demonstrated their capacity as intermediary agents with more nuanced roles, facilitating the accomplishment of myriad tasks with unprecedented efficiency and adaptability (Wang et al. 2024; Xi et al. 2023; Zeng et al. 2023). The Agent paradigm has been extensively applied across multiple domains and tasks, playing a profound role (Qian et al. 2023a, b; Hong et al. 2023). The core of this paradigm lies in introducing external information to change the input distribution of the model, thereby enhancing the model’s performance.
Figure 1: In OOD detection, the Vector Triangle Relationship alters the traditional Binary Relationship by introducing Agents, thereby more effectively processing and distinguishing between ID data and OOD data.
An important and challenging task within the field of machine learning is to enhance the robustness of models across diverse scenarios. When an artificial intelligence system encounters data that significantly deviates from its training data distribution, OOD detection becomes crucial for ensuring its reliability and robustness. In the past, most OOD detection methods employed single-modal learning (Yang et al. 2021; Liu et al. 2021; Hendrycks and Gimpel 2016). As CLIP (Radford et al. 2021) has demonstrated astonishing performance across various downstream tasks, an increasing number of CLIP-based methods for out-of-distribution (OOD) detection have emerged (Ramesh et al. 2022; Wang et al. 2022a; Crowson et al. 2022; Wang, Xing, and Liu 2021; Gao et al. 2024).
@@
In the realm of OOD detection, the predominant methodology consists of binary segmentation between ID and OOD data, typically facilitated by scoring functions. The effectiveness of this prototype is closely tied to the sophisticated design of these scoring functions. Inspired by the remarkable success of the agent-based paradigm, introducing external information as agents in CLIP-based OOD detection framework could reshape the distribution of ID and OOD inputs. Structurally, this paradigm shift from binary segmentation to a vector triangle relationship relational framework holds substantial potential for uncovering deeper insights, potentially transforming the interplay between ID labels and data inputs (Figure  1).
@@
In this paper, we propose the Concept Matching with Agent (CMA) methodology, which integrates neutral textual concept prompts in natural language as Agents within the CLIP-based OOD detection framework. Our framework is illustrated in Figure 2. These agents serve dual roles: as observers and as intermediate hubs that facilitate the interaction between ID Labels and Data inputs. By doing so, we aim to establish a vector triangle relationship among the ID labels, data inputs, and the agents themselves. In this triangular vector relationship (Figure  1), the score of OOD data is diminished due to the collision effect of the Agents, thereby widening the gap between the scores of ID data and OOD data. In other words, images closer to the ID class are less likely to be influenced by neutral text concepts, whereas images from the OOD class are more susceptible to these concepts, resulting in lower scores. The method to achieve our idea is derived from our profound insights into the Language-Vision representation (See Section 3). Building on these insights, we formulate the entire Concept Matching with Agent (CMA) framework.
In summary, our proposed method, CMA, possesses three distinct advantages. (1) Our approach obviates the need for training data, enabling zero-shot OOD detection while leveraging the collision between Agent and both ID and OOD data to effectively widen the gap between the two, ensuring robust performance. This stands in stark contrast to traditional OOD detection methods, which rely on extensive external data for intricate training (Yang et al. 2021; Liu et al. 2021). (2) Our method exhibits remarkable scalability, allowing for the tailoring of specialized Agents to suit various scenarios, thereby further enhancing performance (See Section 5). This is facilitated by the flexibility of our triangular vector relationship, which can enhance the impact of certain OOD images through specific Agents, thereby reducing their scores. (3) It is noteworthy that the CMA maintains robustness against both hard OOD inputs, encompassing both semantic hard OODs (Winkens et al. 2020) and spurious OODs (Ming, Yin, and Li 2022). This makes our approach a truly practical and viable option. The contributions of our study are summarized as follows:
@@ 2 Preliminaries
Contrastive Vision-Language Models.
In comparison to traditional CNN architectures, the ViT (Dosovitskiy et al. 2020) leverages the Transformer Encoder framework (Vaswani et al. 2017) to accomplish the task of image classification, realizing the possibility of utilizing language model architectures for visual tasks. This provides insights into the field of vision-language representation learning, with CLIP (Radford et al. 2021) being a notable representative. CLIP employs self-supervised contrastive objectives to embed images and their corresponding textual descriptions into a shared feature space, achieving alignment between the two. Structurally, CLIP, which utilizes a dual-stream architecture, comprises an image encoder $\mathcal{I}:x\to\mathbb{R}^{d}$ and a text encoder $\mathcal{T}:t\to\mathbb{R}^{d}$. After pretraining on a dataset of 400 million text-image pairs, the joint visual-language embedding of CLIP associates objects in various patterns. Due to the robust performance of CLIP, several OOD detection methods based on it have emerged. Nevertheless, the challenge of how to better utilize Language-Vision representation for OOD detection remains a difficult yet significant issue.
Zero-shot OOD Detection.
For traditional OOD detection frameworks (Hendrycks and Gimpel 2016), a common assumption is made of a typical real-world scenario wherein classifier $f$ are trained on ID data categorized as $\mathcal{Y}_{in}=\left\{1,...,C\right\}$ and subsequently deployed in an environment containing samples from unknown classes $y\notin\mathcal{Y}_{in}$, outside the distribution of the ID. The classifier $f$ is then tasked with determining membership to the ID. In contrast, zero-shot OOD detection (Ming et al. 2022a; Miyai et al. 2023; Fort, Ren, and Lakshminarayanan 2021), in line with the current trend of deep learning, leverages pre-trained models on open datasets, eliminating the need for additional training of the model. The approach determines membership to the ID by calculating results through mapping the data into a common space $\mathbb{R}^{d}$. Compared to traditional OOD detection frameworks that necessitate training, the zero-shot OOD detection framework is notably more versatile and practical.
CLIP-based OOD Detection
For current CLIP-based OOD detection, MCM (Ming et al. 2022a) has become a basic paradigm. Its core idea is to treat text embeddings as “concept prototypes” and evaluate their in-distribution or out-of-distribution properties by measuring the similarity between the input image features and these concept prototypes. Specifically, given a set of ID categories $\mathcal{Y}_{in}$ with a corresponding text description for each category, we first use a pre-trained text encoder $\mathcal{T}$ to convert these text descriptions into $d$ dimensional vectors $\mathbf{c}_{i}=\mathcal{T}(t_{i})\in\mathbb{R}^{d}$, where $i\in\{1,2,...,N\}$ and $N$ is the number of ID categories. For any input image $\mathbf{x}^{\prime}$ whose visual features are extracted by an image encoder $\mathcal{I}$ as $\mathbf{v}^{\prime}=\mathcal{I}(\mathbf{x}^{\prime})\in\mathbb{R}^{d}$, the MCM score is defined as the cosine similarity between the visual features and the closest concept prototype, which is scaled by an appropriate softmax to enhance the separability of ID and OOD samples:
where $\text{sim}(\cdot,\cdot)$ denotes cosine similarity, $\hat{y}=\arg\max_{i}\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{i})$ denotes the most matching concept index, and $\tau$ is a temperature parameter used to adjust the distribution of similarity.
@@ 3 Method
3.1 Out-of-Distribution Detection
Out-of-Distribution (OOD) detection pertains to the task of discerning whether a given input sample originates from the same distribution as the training data (in-distribution, ID) or from a different distribution (out-of-distribution, OOD). Formally, let $X$ denote the input space, with $P_{\text{in}}$ denoting the probability distribution of ID data and $P_{\text{out}}$ denoting the distribution of OOD data. The objective of OOD detection is to classify an input sample $x\in X$ as either belonging to $P_{\text{in}}$ or $P_{\text{out}}$. This problem can be framed as a binary classification challenge, where the model output, denoted as $f(x;\theta)$, yields a probability score $p$ indicating the confidence that $x$ is in-distribution: $p(x)=P(y=1|x;\theta)$, with $y=1$ signifying that $x$ is in-distribution. A threshold $\tau$ is established to facilitate classification, whereby if $p(x)>\lambda$, the sample is classified as in-distribution, and if $p(x)\leq\lambda$, it is classified as out-of-distribution. Thus, OOD detection aims to enhance the robustness and reliability of machine learning models by effectively mitigating the risks posed by unseen or anomalous data.
@@ 3.2 How to construct vector triangle relationships using Agents
Figure 3: Heatmaps depicting the cosine similarity between image inputs and ID concept vectors. In Figure (A), the ID concept vectors consist of sentences containing the word ”cat” of varying lengths. It is observed that images tend to align with longer sentences regardless of whether there is a matching ID concept. Concurrently, keywords such as ”white” significantly influence image matching. In Figure (B), aside from ”cat”, no ID concept can be precisely matched with the given images. However, other than cat images, all images exhibit a preference for aligning with a long sentence devoid of tangible objects. Notably, cat image remains unaffected, aligning solely with the ID concept ”cat”. All the data in the figure is obtained from the practical use of CLIP (https://github.com/openai/CLIP).
@@
    Let $L$ denote the length of the prompt, and $S(L)$ denote the matching score. This relationship could be described as follows:
    $S(L)\propto L\quad\text{for }L\in[a,b],$
    where $a$ and $b$ denote the lower and upper bounds of the prompt length, respectively.
  • Different words have different weights in the text description. The more important the word is in describing the overall image, the higher its weight. This means that keywords such as color and shape have a significant impact on the matching score.
    Let $w_{i}$ denote the weight of the $i$-th word, and let $P$ denote the set of all words in the textual description. The overall weight of the description can be expressed as follows:
    $W(P)=\sum_{i=1}^{n}w_{i}\cdot f_{i}(P)$
    where $f_{i}(P)$ is a function associated with the $i$-th word, encompassing attributes such as word frequency, significance, and so forth, and $n$ denotes the total number of words in the description.
@@
    Let $ID$ denote the set of images in the ID category, $OOD$ denote the set of images in the OOD category, $N$ denote the set of neutral prompts, and $\Delta S(i,p)$ denote the change in score for an image $i$ when a prompt $p$ is applied. The following relationships hold:
    $\forall i\in ID,\forall n\in N:|\Delta S(i,n)|\approx 0$
    $\forall i\in OOD,\forall n\in N:\Delta S(i,n)\ll 0$
    where $|\Delta S(i,n)|\approx 0$ indicates that the change in score for images in the ID category remains approximately zero when neutral prompts are applied, and $\Delta S(i,n)\ll 0$ indicates that the change in score for images in the OOD category is significantly negative when neutral prompts are applied.
The causes of these phenomena can be contemplated through the training methods and mechanisms of the CLIP-like models.Given that the common text input for pre-trained models during training is long sentences, they possess higher confidence in inferring long sentences. Additionally, CLIP-liked models undergo contrastive learning training, which enables the model to autonomously focus on the differences between various text descriptions, thereby giving higher weights to key words and skewing the predicted outcomes towards the positive class without being influenced by other words. We conducted a more detailed discussion and statistical verification in the Appendix.
Drawing on these significant observations and analyses, we propose the inclusion of neutral prompts unrelated to the ID category as Agents. By engaging Agents in a ‘collision’ with OOD and ID data, we aim to distance the OOD data from the ID domain.
@@ 3.3 Proposed approach
Based on the aforementioned analysis, Concept Matching with Agent (CMA) employs neutral prompts as Agents to widen the gap between ID and OOD by constructing a vector triangle relationship. Essentially, CMA employs external agents to minimize the maximum score, which is fundamentally distinct from traditional learning methods. Therefore, it does not require additional data or training. Specifically, for a set of textual descriptions $t_{i}$, where $i\in\{1,2,...,N\}$ corresponding to ID categories $\mathcal{Y}_{in}$, we select a certain number of neutral prompts $\mathcal{Y}_{ntc}$ to concatenate with them. The number is generally the same as the number of ID categories $N$, and the concatenated textual concepts $\mathcal{Y}_{con}$ are used as the input of the text encoder $\mathcal{T}$ to output text features $\mathbf{c}_{i}=\mathcal{T}(t_{i})$, where $i\in\{1,2,...,2N\}$, $t_{i}\in\mathcal{Y}_{con}=\{\mathcal{Y}_{in},\mathcal{Y}_{ntc}\}$. Then, it is calculated with the image features output $\mathbf{v}^{\prime}=\mathcal{I}(\mathbf{x}^{\prime})$ by the image encoder $\mathcal{I}$.
@@
where $\text{sim}(\cdot,\cdot)$ denotes cosine similarity, $\hat{y}=\mathop{\arg\max}\limits_{1\leq i\leq N}\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{i})$ denotes the most matching concept index in ID categories, and $\tau$ is a temperature parameter used to adjust the distribution of similarity.
@@
$g\left(\mathbf{x}^{\prime};\mathcal{Y}_{\mathrm{in}},\mathcal{Y}_{ntc},\mathcal{T},\mathcal{I}\right)=\left\{\begin{array}[]{ll}1,&S_{\mathrm{CMA}}\geq\lambda\\ 0,&otherwise\end{array},\right.$